| Operation Summary for February 27, 2012 - February 29, 2012 |
|
Monday February 27 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam. During the day shift, a Booster RF station (BRF3) tripped off at 8:09 AM; it reset. At 8:09 AM, BRF11 tripped off for the first of three times in four hours; it reset. Operations reported problems with the Booster RF system sparking and, at 8:16 AM, a RF technician investigated BRF16. The NuMI beam permit tripped off at 8:31 AM; it reset. At 8:41 AM, BRF14 tripped off; it reset. The NuMI beam permit tripped off at 8:543 AM due to an MI incomplete RF gap indication, but with no MI RF station trips; Operators reset the permit. At 9:02 AM, BRF19 tripped off; it reset. Water Group personnel reported at 9:28 AM that they were adding water to the MI RF LCW system. A Booster person began a two-hour study at 10:01 AM. At 10:20 AM, the MI RF South Anode Power Supply tripped off; it reset. At 11:09 AM, a Linac RF station (LRF3) tripped off; it reset. At 1:18 PM, BRF19 tripped off; it reset. A Booster person investigated the RF sparking problems at 1:22 PM. At 1:24 PM, BRF3 tripped off; it reset. A MI person began a MI aperture scan at 3:01 PM. Operators had completed a search and secure Neutrino Area enclosures N01 through to NM3 at 3:41 PM. During the evening shift, Operations received permission to send beam to the FTBF Meson experiment T-1015 at 5:07 PM. Operations established beam to T-1015 at 5:42 PM and disabled their beam at 7:49 PM. At 10:05 PM, the MI South Anode Power Supply tripped off; it reset. Operations completed a Booster injection tune at 10:39 PM. At 11:16 PM, a Linac quadrupole magnet power supply (L:QPS104) tripped off; it reset. Tuesday February 28 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 1:37 AM, BRF1, 2 and 7 tripped off; they all reset. At 5:05 AM, a Linac Klystron RF station (KRF1) tripped off; it reset. Operations completed a 750 KeV line tune. MI personnel began a MI/ Recycler study at 7:08 AM. At 7:21 AM, the NuMI beam permit tripped off; it reset. At 7:26 AM, Neutron Therapy Facility personnel prepared to treat three patents later in the morning. A Booster person began a study at 7:46 AM. During the day shift, KRF5 tripped off at 9:02 AM; it reset. A Switchyard person began a Meson MTest/MCenter beam-splitting study at 10:08 AM. At 12:58 PM, BRF14 tripped off; it reset. Operations established beam to MTA around 1:28 PM. During the evening shift, at 4:10 PM, Operations finished a search and secure for N01 and NM2 and turned on all associated power supplies. Operations disabled beam to MTA at 5:11 PM. At 9:17 PM, LRF2 tripped off for the first of several trips. Operators ended up increasing the driver screen voltage and a Linac expert came in to tune on the station. Operations disabled beam to MTest at 9:18 PM. Wednesday February 29 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, Operations monitored NuMI and MiniBooNE. Operations Coordinator (OpsCo) Plans for Today The OpsCo will schedule a short downtime to access the H- dome and to reboot the LLRF for MI/RR. Continue to deliver beam to users. |
|
More Information Comments and Suggestions |
| last modified 02/29/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |